HOW WOULD I USE SYNTHETIC DIVISION GIVEN THE POLYNOMIAL:
P(X)=X^4-4X^3-6X^2-4X-15 AND THE FACTOR THEOREM TO FIND WHETHER X-(1+2i) IS A FACTOR
THANKS TANYA

Answer by AnlytcPhil(1116)

HOW WOULD I USE SYNTHETIC DIVISION GIVEN THE POLYNOMIAL:
P(X)=X^4-4X^3-6X^2-4X-15 AND THE FACTOR THEOREM TO FIND WHETHER
X-(1+2i) IS A FACTOR

Actually it's not as I will show you, but first it will be more 
instructive to you if I first demonstrate a case of a similar
polynomial for which x-(1+2i) is a factor:

Find whether x-(1+2i) is a factor of 

P(x) = x4 - 6x3 + 18x2 - 30x + 25

We start out with the synthetic division algorithm. Start as 
usual by bringing down the 1:

1+2i| 1  -6     18     -30      25
    |                                
      1                             

Multiply the 1 by 1+2i and put it diagonally above the 1 under 
the -6

1(1+2i) = 1+2i


1+2i| 1  -6     18     -30      25
    |     1+2i                                         
      1                             

Add -6 and 1+2i, getting -5+2i, and write that on the bottom line


1+2i| 1  -6     18     -30      25
    |     1+2i                                         
      1  -5+2i                    

Multiply the (-5+2i) by (1+2i)

(-5+2i)(1+2i) = -5-10i+2i+4iČ = -5-8i+4(-1) = -5-8i-4 = -9-8i

Write that diagonally above -5+2i

1+2i| 1  -6     18     -30      25
    |     1+2i  -9-8i                        
      1  -5+2i                     

Add 18 and -9-8i, getting 9-8i and write that on the bottom 
line


1+2i| 1  -6     18     -30      25
    |     1+2i  -9-8i                          
      1  -5+2i   9-8i              

Multiply 9-8i by 1+2i

(9-8i)(1+2i) = 9+18i-8i-16iČ = 9+10i-16(-1) = 9+10i+16 = 25+10i

Write that diagonally above 9-8i, then add it to -30 getting 
-5+10i

1+2i| 1  -6     18     -30      25
    |     1+2i  -9-8i   25+10i         
      1  -5+2i   9-8i   -5+10i            

Multiply -5+10i by 1+2i

(-5+10i)(1+2i) = -5-10i+10i+20iČ = -5+20(-1) = -5-20 = -25

Write that diagonally above -5+25, add to 25 and get 0 remainder


1+2i| 1  -6     18     -30      25
    |     1+2i  -9-8i   25+10i -25
      1  -5+2i   9-8i   -5+10i   0

So we see that since we got a 0 remainder we can say that

x - (1+2i) IS a factor of x4 - 6x3 + 18x2 - 30x + 25.

========================================================

Now if we do your problem the same way, we get

1+2i| 1  -4     -6     -4     -15
    |     1+2i  -7-4i  -5-30i  51-48i
      1  -3+2i -13-4i  -9-30i  36-48i

So as you see we do not get a 0 remainder, so NO, x-(1+2i)
is NOT a factor of P(x) = x4 - 4x3 - 6x2 - 4x - 15.
as it was in the example I gave.  


Edwin